OpenStudy (anonymous):

Solve |2 - 3x| = 7. A. {-5/3, 3} B. {-15, 3} C. {-3, 5/3} D. {-3, 15}

OpenStudy (kameron018):

A

OpenStudy (aaronq):

When you have an absolute sign, you need to take into account the negative and positive possibilities. |2 - 3x| = 7 2 - 3x = 7 -(2 - 3x)= 7 x = -5/3 x = 3

OpenStudy (cj49):

A